Circular No. 5981

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J1948+32 = GRO J2014+34 D. Chakrabarty and T. A. Prince, California Institute of Technol-

ogy; M. H. Finger, Universities Space Research Association; R. B. Wilson, Marshall Space Flight Center, NASA; and G. N. Pendleton, University of Alabama, Huntsville, report for the Compton Gamma Ray Observatory BATSE team: “Further observations of the recently-discovered, 18.7-s, transient hard x-ray pulsar in Cygnus (originally designated GRO J2014+34; see IAUC 5977) have yielded a more precise position. The estimated 90-percent-confidence error box is approximately a 2^{o}.5 × 5^{o}.5 quadrilateral, with corners at α = 19^{h}58^{m}, δ = +29^{o}.0; 19^{h}50^{m}, +34^{o}.0; 19^{h}34^{m}, +35^{o}.0; 19^{h}46^{m}, +28^{o}.0 (equinox 2000.0). In accordance with the substantially improved position, we have redesignated the source as GRO J1948+32. The pulsed emission was first detectable on Apr. 6 UT and continues to be visible as of Apr. 28. Optical and soft-x-ray observations of the error box are strongly encouraged in order to further constrain the x-ray position and identify the accreting companion.”

SUPERNOVA 1994I IN NGC 5194 R. Kirshner, Center for Astrophysics, reports on behalf of the Su-

pernova INtensive Studies (SINS) team that SN 1994I was observed with the Hubble Space Telescope (HST) Faint Object Spectrograph (FOS) and Planetary Camera on Apr. 18 and 19 UT. The spectrum extends from 160 to 470 nm; it is consistent with the type-Ic classification of earlier optical data, but the ultraviolet shows a strong resemblance to the FOS spectrum of the type-Ia SN 1992A about 5 days after maximum light (Kirshner et al. 1993, Ap.J. 415, 589). The spectrum shows no signature of the circumstellar matter inferred from radio observations. Ultraviolet flux levels from the spectrum and from U-band images are consistent at ∼ 8 × 10^{−16} erg s^{−1} cm^{−2} A^{−1}, and are much lower at shorter wavelengths. The supernova appears projected on diffuse emission, possibly from a spiral arm in M51. Comparison of the recent U image with 1992 HST observations of M51 demonstrates that the star suggested by Richmond (IAUC 5966) as the possible progenitor is present both before and after the eruption of SN 1994I, ∼ 0′′.36 from the supernova site. The absence of an identified progenitor in the earlier HST data may provide useful constraints. Further HST images and spectra are planned for May 12 and 14.”

1994 April 29 Daniel W. E. Green